plaease help me use the substitution method to solve the linear system.
y = x-4
4x+y = 26

Answer by ilana(307)

By "substitution," we mean starting with 2 equations and 2 variables, and replacing one variable to end up with 1 equation and 1 variable, which we then solve.
Here, we have an equation of the form y=(something in terms of x), so it is simplest to substitute that equation for y in the other equation.
So, we take x-4 for y and use that in the equation 4x+y=26, giving us 4x+(x-4)=26.
We then simplify this to get 4x+x-4=26, so 5x-4=26, so 5x=30, and finally x=6.
We then need only the value of y, so we use x=6 and y=x-4 to get y=(6)-4=2.
So x=6 and y=2.
We can check these values in both equations and they should make the statements true:
y=x-4: 2=6-4=2, so this is true.
4x+y=26: 4(6)+(2) = 26
24+2=26 = 26, so this is true as well.